OpenStudy (anonymous):

I need some information on radio waves. Will you guys give me some, please? MEDAL FOR THE BEST ANSWER!!!! (:

OpenStudy (anonymous):

I dont know much but what do you need to know?

OpenStudy (anonymous):

Well. me and my friends are doing a project on the electromagnetic spectrum and i need to do research on radio waves, all i need is half a page of research so just tell me what you know.

OpenStudy (anonymous):

radio waves are electromagnetic wave (same as light) that goes on the scale of 3khz - 300ghz. as all EM waves they follow the waves equation: \[f(x,t) = y(x-vt)+h(x+vt)\] or\[y=Asin[k(x-vt+\phi)] = Asin(kx-omega t+\phi)\] where A is the amplitude of the wave. k is const. v is the speed t is the time phi is the place from which the wave starts and omega is the frequency. the radio you have at home usually uses 2 types of radio waves transmission: AM and FM. in order to send the information the radio station encodes it on another wave called Carrier signal. the AM coding uses the difference in Amplitude in order to "form" the information on the carrier signal and the FM coding uses the difference in Frequency in order to do so. that's most of it. I can give you some more info about EM waves in general, but if this is for high school than this info will be over your Math ability of understanding. other than that you can always look on radio waves in wikipedia I think you can get half a page from there

OpenStudy (anonymous):

OMG. Thank you SO much!!!!!

OpenStudy (anonymous):

np :)